40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Dec 10 17:03:41 2019, anchal, DailyProgress, Laser, Laser Settings back to defaults 
    Reply  Wed Dec 18 14:04:43 2019, anchal, DailyProgress, Laser, South Laser Power Dropped. SouthLaserPowerDropped.pdf
Message ID: 2492     Entry time: Tue Dec 10 17:03:41 2019     Reply to this: 2501
Author: anchal 
Type: DailyProgress 
Category: Laser 
Subject: Laser Settings back to defaults 

I put laser settings on both North and South Cavities back to default. From this point onwards, all settings about the lasers would be known and kept track of. The red values are the settings that were changed.

NPRO Laser Settings

Property Display Symbol North South Units Notes
Laser Model - M126N-1064-700, SN 5519, Dec 2006 126N-1064-500, SN 280, Nov 1997 -  
Diode Temperature DT 22.3 28.7 ^\circ C Informational only.
Diode TEC Voltage DTEC 0.8 0.7 V Informational only. +ve -> cooling, -ve -> heating.
Measured Laser Crystal Temperature LT 40.8 55.2 ^\circ C Informational only. Calibration dependent.
Laser TEC Voltage LTEC 0.0 -0.5 V Informational only. +ve -> cooling, -ve -> heating. Manual says typically should be 0.0V.
Target Laser Crystal Temperature T 40.087 -> 40.0010 48.0010 ^\circ C Changed back to factory set value on North Side.
Laser Head Power Level PWR 66  ->  624 92  ->  101 mW Calibration dependent. CHanged the calibration to meet the power meter but even then, power meter says a maximum 500 mW, so North Side is not entirely correct. On the south side, it was difficult to mount power meter perpendicular to the beam, so there might be some clipping loss in calibration.
Power Adjustment ADJ 0 -2  ->  0 - From -50 (off) to +10. Changed the diode current around set value.
Diode Current DC 2.06 2.04 A It can be changed to change power level. Reflects measured value.
Diode Power Monitor DPM 0.00 0.00 V

Calibration Dependent.

Noise Easter NE ON ON - -
Laser Diode Status LD ON ON -  
Nominal Diode Current - All the way clockwise All the way clockwise - It can be changed by turning the left potentiometer from the back of the laser head. Factory default is all the way clockwise. I have set both North and South Lasers to this point.

Note:

While turning the nominal diode current of south laser all the way clockwise, I found that the laser power peaks before the maximum diode current is reached. This diode current is about 1.9 A. This is unexpected. Any explanations on this would be helpful.

ELOG V3.1.3-